### v3

@@ 4.2 Propositions and Lemmas of $\sigma_{i}$ neural networks
(12) $\widetilde{g}(x)=\sum_{i=0}^{\lfloor N^{1/d}\rfloor-1}s\left(4Kx-4i-\frac{1}{2}\right)$
@@
Next, we construct $\psi_{i}$ for $i=2,\ldots,\lceil\frac{2}{d}\log_{2}L\rceil+2$ based on the symmetry and periodicity of $\widetilde{g}(x)$ due to the symmetry of $s(x)$ (see Fig. 2). $\psi_{2}$ is the function with period $\frac{2}{\lfloor N^{1/d}\rfloor\lfloor L^{2/d}\rfloor}$ in $\left[0,\frac{1}{\lfloor L^{2/d}\rfloor}\right]$, and each period is a hat function with gradient 1. $\psi_{i}$ for $i=3,\ldots,\lceil\frac{2}{d}\log_{2}L\rceil+2$ is the hat function in $\left[0,\frac{2^{i-2}}{\lfloor L^{2/d}\rfloor}\right]$. The schematic diagram is in Figure 3 (The diagram is shown the case for $\lfloor N^{1/d}\rfloor$ is a even integer).
@@
Note that $\psi_{2}\circ\psi_{3}\ldots\circ\psi_{\lceil\frac{2}{d}\log_{2}L\rceil+2}(x)$ is the function with period $\frac{2}{\lfloor N^{1/d}\rfloor\lfloor L^{2/d}\rfloor}$ in $[0,1]\subset\left[0,\frac{2^{\lceil\frac{2}{d}\log_{2}L\rceil}}{\lfloor L^{2/d}\rfloor}\right]$, and each period is a hat function with gradient 1. Then function
(13) $\psi(x):=\widetilde{g}(x)\circ\psi_{2}\circ\psi_{3}\ldots\circ\psi_{\lceil\frac{2}{d}\log_{2}L\rceil+2}(x):=\widetilde{g}\circ\psi_{1}$
@@
$\psi_{2}$ is a network with width $4\lfloor N^{1/d}\rfloor$ and one hidden layer. $\psi_{i}$ for $i=3,\ldots,\lceil\frac{2}{d}\log_{2}L\rceil+2$ is a $\sigma_{1}$-NN with width 3 and one hidden layer. Hence $\psi_{1}(x)$ is a $\sigma_{1}$-NN with width $4\lfloor N^{1/d}\rfloor$ and depth $\lceil\frac{2}{d}\log_{2}L\rceil$ and $s_{1}(x)=\psi\left(x+\frac{1}{8K}\right)$ and $s_{2}(x)=\psi\left(x+\frac{5}{8K}\right)$.
Based on (ii) in Lemma 4.3, we have that there is a $\sigma_{1}$ neural network $\lambda_{\boldsymbol{m},1}(\boldsymbol{x})$ with the width $4N$ and depth $\lceil 2\log_{2}L\rceil$ and a $\sigma_{2}$ neural network $\lambda_{\boldsymbol{m},2}(\boldsymbol{x})$ with the width $6N$ and depth $\lceil 1+\log_{2}d\rceil$, such as
$\lambda_{\boldsymbol{m},2}\circ\lambda_{\boldsymbol{m},1}(\boldsymbol{x})=\prod_{j=1}^{d}s_{m_{j}}\left(\frac{x_{j}-a}{b-a}\right)=s_{\boldsymbol{m}}(\boldsymbol{x}),\boldsymbol{x}\in[a,b]^{d},$
where the order of width is based on $dN^{1/d}\leq N$ for $N>d$.
@@
Proposition 4.10 ([26]).
Given any $N,L\in\mathbb{N}_{+}$ and $\delta\in\Big{(}0,\frac{1}{3K}\Big{]}$ for $K=\lfloor N^{1/d}\rfloor^{2}\lfloor L^{2/d}\rfloor$, there exists a $\sigma_{1}$-NN $\phi$ with the width $54N+$ 5 and depth $44L+$ 4 such that
$\phi(x)=k,x\in\left[a+\frac{k\cdot(b-a)}{K},a+\frac{(k+1)\cdot(b-a)}{K}-\delta\cdot 1_{k<a+(K-1)\cdot(b-a)}\right],~{}k=0,\ldots,K-1.$
Remark 4.11.
@@
4.3.1 Function approximation in Sobolev spaces through the Bramble–Hilbert Lemma [4, Lemma 4.3.8]
Before we show Proposition 4.13, we define subsets of $\Omega_{\boldsymbol{m}}$ for simplicity notations.
For any $\boldsymbol{m}\in\{1,2\}^{d}$, we define
(14) $\Omega_{\boldsymbol{m},\boldsymbol{i}}:=[a,b]^{d}\cap\prod_{j=1}^{d}\left[a+\frac{2i_{j}-1_{m_{j}<2}}{2K}\cdot(b-a),a+\frac{3+4i_{j}-2\cdot 1_{m_{j}<2}}{4K}\cdot(b-a)\right]$
$\boldsymbol{i}=(i_{1},i_{2},\ldots,i_{d})\in\{0,1\ldots,K\}^{d}$, and it is easy to check $\bigcup_{\boldsymbol{i}\in\{0,1\ldots,K\}^{d}}\Omega_{\boldsymbol{m},\boldsymbol{i}}=\Omega_{\boldsymbol{m}}$.
Proposition 4.13.
@@
Let $K\in\mathbb{N}_{+}$, $p\in[1,\infty]$ and $n\geq 2$. Then for any $f\in W^{n,p}([a,b]^{d})$ and $\boldsymbol{m}\in\{1,2\}^{d}$, there exist piece-wise polynomials function $f_{K,\boldsymbol{m}}=\sum_{|\boldsymbol{\alpha}|\leq n-1}g_{f,\boldsymbol{\alpha},\boldsymbol{m}}(\boldsymbol{x})\boldsymbol{x}^{\boldsymbol{\alpha}}$ on $\Omega_{\boldsymbol{m}}$ (Definition 4.5) with the following properties:
$\|f-f_{K,\boldsymbol{m}}\|_{W^{2,p}(\Omega_{\boldsymbol{m}})}$ $\leq C_{1}(n,d)\|f\|_{W^{n,p}(\Omega_{\boldsymbol{m}})}K^{-(n-2)},$
$\|f-f_{K,\boldsymbol{m}}\|_{W^{,p}(\Omega_{\boldsymbol{m}})}$ $\leq C_{1}(n,d)\|f\|_{W^{n,p}(\Omega_{\boldsymbol{m}})}K^{-(n-1)},$
(15) $\|f-f_{K,\boldsymbol{m}}\|_{L^{p}(\Omega_{\boldsymbol{m}})}$ $\leq C_{1}(n,d)\|f\|_{W^{n,p}(\Omega_{\boldsymbol{m}})}K^{-n}.$

### v4

@@ 4.2 Propositions and Lemmas of $\sigma_{i}$ neural networks
$\widetilde{g}(x)=\sum_{i=0}^{\lfloor N^{1/d}\rfloor-1}s\left(4Kx-4i-\frac{1}{2}\right)$ (4.10)
@@
Next, we construct $\psi_{i}$ for $i=2,\ldots,\lceil\frac{2}{d}\log_{2}L\rceil+2$ based on the symmetry and periodicity of $\widetilde{g}(x)$ due to the symmetry of $s(x)$. $\psi_{2}$ is the function with period $\frac{2}{\lfloor N^{1/d}\rfloor\lfloor L^{2/d}\rfloor}$ in $\left[0,\frac{1}{\lfloor L^{2/d}\rfloor}\right]$, and each period is a hat function with gradient 1. $\psi_{i}$ for $i=3,\ldots,\lceil\frac{2}{d}\log_{2}L\rceil+2$ is the hat function in $\left[0,\frac{2^{i-2}}{\lfloor L^{2/d}\rfloor}\right]$. The schematic diagram is in Fig. 4 (The diagram is shown the case for $\lfloor N^{1/d}\rfloor$ is a even integer).
@@
Note that $\psi_{2}\circ\psi_{3}\ldots\circ\psi_{\lceil\frac{2}{d}\log_{2}L\rceil+2}(x)$ is the function with period $\frac{2}{\lfloor N^{1/d}\rfloor\lfloor L^{2/d}\rfloor}$ in $0,1]\subset\left[0,\frac{2^{\lceil\frac{2}{d}\log_{2}L\rceil}}{\lfloor L^{2/d}\rfloor}\right]$, and each period is a hat function with gradient 1. Then function
$\psi(x):=\widetilde{g}(x)\circ\psi_{2}\circ\psi_{3}\ldots\circ\psi_{\lceil\frac{2}{d}\log_{2}L\rceil+2}(x):=\widetilde{g}\circ\psi_{1}$ (4.11)
@@
$\psi_{2}$ is a network with width $4\lfloor N^{1/d}\rfloor$ and one hidden layer. Rest $\psi_{i}$ is a $\sigma_{1}$-NN with width 3 and one hidden layer. Hence $\psi_{1}(x)$ is a $\sigma_{1}$-NN with width $4\lfloor N^{1/d}\rfloor$ and depth $\lceil\frac{2}{d}\log_{2}L\rceil$ and $s_{1}(x)=\psi\left(x+\frac{1}{8K}\right)$ and $s_{2}(x)=\psi\left(x+\frac{5}{8K}\right)$.
Based on (ii) in Lemma 3, we have that there is a $\sigma_{1}$ neural network $\lambda_{\bm{m},1}(\bm{x})$ with the width $4N$ and depth $\lceil 2\log_{2}L\rceil$ and a $\sigma_{2}$ neural network $\lambda_{\bm{m},2}(\bm{x})$ with the width $6N$ and depth $\lceil 1+\log_{2}d\rceil$, such as
$\lambda_{\bm{m},2}\circ\lambda_{\bm{m},1}(\bm{x})=\prod_{j=1}^{d}s_{m_{j}}\left(\frac{x_{j}-a}{b-a}\right)=s_{\bm{m}}(\bm{x}),\bm{x}\in[a,b]^{d},$
where the order of width is based on $dN^{1/d}\leq N$ for $N>d$. ∎
@@
Proposition 5 ([39]).
Given any $N,L\in\mathbb{N}_{+}$ and $\delta\in\Big{(}0,\frac{1}{3K}\Big{]}$ for $K=\lfloor N^{1/d}\rfloor^{2}\lfloor L^{2/d}\rfloor$, there exists a $\sigma_{1}$-NN $\phi$ with the width $4N+5$ and depth $4L+4$ such that
$\phi(x)=k,x\in\left[a+\frac{k\cdot(b-a)}{K},a+\frac{(k+1)\cdot(b-a)}{K}-\delta\cdot 1_{k<a+(K-1)\cdot(b-a)}\right],$
for $k=0,1,\dots,K-1.$
Remark 2.
@@ 4.3 Proof of Theorem 2
4.3.1 Proof of Theorem 2
Define subsets of $\Omega_{\bm{m}}$ for simplicity notations.
For any $\bm{m}\in\{1,2\}^{d}$, we define
$\Omega_{\bm{m},\bm{i}}:=[a,b]^{d}\cap\prod_{j=1}^{d}\left[a+\frac{2i_{j}-1_{m_{j}<2}}{2K}\cdot(b-a),a+\frac{3+4i_{j}-2\cdot 1_{m_{j}<2}}{4K}\cdot(b-a)\right]$ (4.12)
$\bm{i}=(i_{1},i_{2},\ldots,i_{d})\in\{0,1\ldots,K\}^{d}$, and it is easy to check $\bigcup_{\bm{i}\in\{0,1\ldots,K\}^{d}}\Omega_{\bm{m},\bm{i}}=\Omega_{\bm{m}}$.
Proposition 7 ([59, Theorem 8]).
@@ 4.3.1 Proof of Theorem 2
Let $K\in\mathbb{N}_{+}$, $p\in[1,\infty]$ and $n\geq 2$. Then for any $f\in W^{n,p}([a,b]^{d})$ and $\bm{m}\in\{1,2\}^{d}$, there exist piece-wise polynomials functions $f_{K,\bm{m}}=\sum_{|\bm{\alpha}|\leq n-1}g_{f,\bm{\alpha},\bm{m}}(\bm{x})\bm{x}^{\bm{\alpha}}$ on $\Omega_{\bm{m}}$ (Definition 3) with the following properties:
$\|f-f_{K,\bm{m}}\|_{W^{2,p}(\Omega_{\bm{m}})}$ $\leq C_{1}(n,d,|b-a|)\|f\|_{W^{n,p}(\Omega_{\bm{m}})}K^{-(n-2)},$
$\|f-f_{K,\bm{m}}\|_{W^{1,p}(\Omega_{\bm{m}})}$ $\leq C_{1}(n,d,|b-a|)\|f\|_{W^{n,p}(\Omega_{\bm{m}})}K^{-(n-1)},$
$\|f-f_{K,\bm{m}}\|_{L^{p}(\Omega_{\bm{m}})}$ $\leq C_{1}(n,d,|b-a|)\|f\|_{W^{n,p}(\Omega_{\bm{m}})}K^{-n}.$ (4.13)